# Proximal Gradient Dynamics: Monotonicity, Exponential Convergence, and Applications

### v1

@@ Abstract
In this letter, we study the proximal gradient dynamics. This recently-proposed continuous-time dynamics solves optimization problems whose cost functions are separable into a nonsmooth convex and a smooth component. First, we show that the cost function decreases monotonically along the trajectories of the proximal gradient dynamics. We then introduce a new condition that guarantees exponential convergence of the cost function to its optimal value, and show that this condition implies the proximal Polyak-Łojasiewicz condition. We also show that the proximal Polyak-Łojasiewicz condition guarantees exponential convergence of the cost function. Moreover, we extend these results to time-varying optimization problems, providing bounds for equilibrium tracking. Finally, we discuss applications of these findings, including the LASSO problem, quadratic optimization with polytopic constraints, and certain matrix based problems.
@@ I Introduction
Optimization problems are central to various domains of engineering and data science. Many of these problems can be framed as minimizing a smooth cost function combined with a nonsmooth but convex regularizer that promotes structure in the solution. For example, the addition of the $\ell_{1}$ norm as a regularizer induces sparsity, a key principle in compressed sensing and sparse coding [3]. In other cases, convex constraints can be incorporated using indicator functions. A common approach for solving such problems, where the cost function is separable into a smooth, possibly nonconvex, and a nonsmooth convex component, is the proximal gradient algorithm [16]. This (discrete-time) algorithm gained prominence in sparse encoding applications [3]. A list of widely used regularizers and their associated proximal operators is available in [2, Chapter 10].
Analyzing the continuous-time versions of discrete-time optimization algorithms has garnered much interest since the seminal work in [1]. Continuous-time dynamics have proven useful across various engineering disciplines, including notable contributions from Hopfield and Tank in dynamical neuroscience [13], Kennedy and Chua in analog circuit design [15], and Brockett in systems and control theory [5]. There are several compelling reasons to investigate continuous-time counterparts of discrete optimization methods. First, continuous-time dynamics can offer insights into the design and convergence properties of their discrete-time equivalents. Second, differential equations that solve optimization problems in continuous time can be implemented on analog circuits [20, 15], which can lead to faster computational solutions. Finally, since biological systems inherently operate in continuous time, designing algorithms in this context opens the door to biologically plausible models that could be implemented via neural circuits. This approach may offer a deeper understanding of brain function [13, 6].
@@
Related Work: The analysis of algorithms through the lens of systems theory is becoming popular [9]. Proximal gradient algorithms have been extensively studied in the discrete-time optimization setting [16]. A key result from this analysis is the identification of the Polyak-Łojasiewicz (PL) condition, initially introduced for gradient descent in [17] and later extended to proximal methods in [14]. When the cost function satisfies this condition in discrete time, proximal gradient algorithms achive linear convergence. The authors of [14] also discuss several existing conditions in literature and their relationships with the PL condition. Notably, they propose proximal versions of the PL and Kurdyka Łojasiewicz (KL) conditions and show their equivalence.
@@
The study of optimization problems as continuous-time dynamical systems has a long history [1, 7, 18, 8]. The continuous-time proximal gradient dynamics were first proposed in [12] where it was shown that the PL condition implies the exponential convergence of the forward-backward envelope, even in the absence of strong convexity.
@@
Contributions: In this work, we analyze the convergence properties and applications of proximal gradient dynamics for optimization problems with cost functions that combine a nonsmooth convex term and a potentially nonconvex smooth term. Our first contribution demonstrates that the cost function decreases monotonically along the trajectories of the proximal gradient dynamics, even in a general nonconvex setting. We then identify a natural condition that ensures exponential convergence in continuous time; we show that our condition implies the proximal KL condition. Moreover, we examine the PL condition for proximal gradient dynamics and prove that it guarantees exponential convergence of the cost function. Additionally, we derive Lipschitz-based bounds for tracking the optimal function value in parameter-varying problems under the PL condition. Finally, we explore applications of these results in sparse reconstruction, quadratic optimization, and the analysis of some matrix based problems.
@@
The structure of the letter is as follows: Section II covers preliminaries and formulates the optimization problem. Section III presents our main result on the convergence of the flow dynamics. The proximal PL condition is discussed in Section IV, followed by applications of our results in Section V.
@@ II-A Proximal operator and the proximal gradient dynamics
The proximal gradient dynamics proposed in [12], are given by,
@@ II-B Polyak-Łojaciewicz and related conditions
A differentiable function $f\colon{}^{n}\rightarrow\real$ which has a minimum at $x^{\star}\in\mathbb{R}^{n}$ is said to satisfy the PL condition if there exists a $\mu>0$ such that,
@@
The PL condition is weaker than the usual assumption of convexity. A detailed discussion of the different conditions for convergence of gradient descent in discrete time is discussed in [14]. Further [14], also proposes an extension of the PL condition for the proximal case, under the assumption of $L-$smoothness.
@@
$f\colon{}^{n}\rightarrow\real$ and $g\colon{}^{n}\rightarrow\overline{\real}$ are said to satisfy the proximal PL condition if there exists a ${\mu>0}$ and ${\alpha>0}$, for all $x$, such that,
@@
Lemma 2 (Equivalence of proximal PL and KL Conditions [14]):
Let $L$ be the smoothness parameter of $f$. The proximal PL condition (7) with $\alpha=1/L$, and the proximal KL Condition (8) with exponent $1/2$ are equivalent to each other, i.e., if a function $F$ satisfies either condition for some parameter value, there is guaranteed to exist a valid parameter value for which the other condition also holds.
@@ III Global convergence of the proximal gradient dynamics
We begin by introducing a technical lemma that provides a closed-form expression for a subgradient of a CCP function evaluated at its proximal operator.
Lemma 3 (Proximal optimality condition):
Let $g\colon{}^{n}\rightarrow\overline{\real}$ be CCP. Then for $x\in{}^{n},\alpha>0$, define $z:=\mathrm{prox}_{\alpha g}(v)$. Then
$\alpha^{-1}\bigl{(}v-z\bigr{)}\in\partial g(z).$ (9)
  •       Proof:
    By definition of $z$ and $\mathrm{prox}$,
    $z$ $=\mathrm{prox}_{\alpha g}(v)=\underset{u\in{}^{n}}{\textup{arg\,min}}\;g(u)+\frac{1}{2\alpha}\|u-v\|_{2}^{2}$
    $\iff 0\in\partial g(z)+\alpha^{-1}(z-v).$
    The result follows. ∎
Next, we present our first main result which shows that under the proximal gradient dynamics, the cost function decreases along trajectories of the system. First, we present a simplified version of the result that assumes $g$ is finite-valued, in order to build intuition and provide some additional results in this special case. Then we extend the theorem to the case of a general, possibly non-finite-valued, $g$.
@@
For the optimization problem (3), let the following assumptions hold true:
@@
    For the proximal gradient dynamics (5), let $\phi_{h}\colon{}^{n}\rightarrow{}^{n}$ denote the $h$-time flow map of the dynamics. Under the assumption of continuously differentiable $f$, $\operatorname{F_{\mathrm{prox}}}$ is locally Lipschitz, and thus the flow map $\phi_{h}$ exists for sufficiently small $h\in\real$ forward in time. We consider the function $V(x)=F(x)-F^{*}$, and consider its Dini derivative with respect to time, denoting $x(t)$ by $x$ for brevity, setting ${z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))}$,
    $D^{+}V(x)$ $=\limsup_{h\to 0^{+}}\frac{V(\phi_{h}(x))-V(x)}{h}$
    $\leq\nabla f(x)^{\top}(z-x)+\limsup_{h\to 0^{+}}\frac{g(\phi_{h}(x))-g(x)}{h}$
    $\leq\nabla f(x)^{\top}(z-x)+\limsup_{h\to 0^{+}}\frac{s_{x}^{\top}(x-\phi_{h}(x))}{h}$
    $=\nabla f(x)^{\top}(z-x)+s_{x}^{\top}(z-x),$ (11)
    for $s_{x}\in\partial g(x)$. By convexity of $g(x)$, we have,
    $D^{+}V(x(t))\leq\nabla f(x)^{\top}(z-x)+g(z)-g(x).$
@@
Under the following assumptions,
  1. (i)
    $f$ is continuously differentiable.
  2. (ii)
    $g$ is CCP, takes finite values for all finite $x$.
The condition (10) implies the KL condition (8), and by extension also implies the PL condition (7).
@@
    The KL condition follows upon squaring both sides and substituting (10), and the PL condition follows due to the equivalence in Lemma 2. ∎
In Theorem 4, we show that costs along the trajectories of the proximal gradient dynamics are decreasing, in cases where $g$ is finite-valued. In many practical examples, for instance, in the case where $g$ is an indicator function, this assumption is not met. However, in Theorem 9, we will show that the cost remains non-increasing even for non-finite-valued $g$
For the proximal gradient dynamics (5), recall that $\phi_{h}\colon{}^{n}\rightarrow{}^{n}$ denotes the $h$-time flow map of the dynamics. We consider the evolution of $F(x)=f(x)+g(x)$ along trajectories of the system by studying the evolution of $t\mapsto F(\phi_{t}(x))$ for all $x$. First, we introduce an auxillary result.
Lemma 8 (Forward-invariance of $\operatorname{Dom}g$):
  •       Proof:
    The proof is a consequence of Nagumo’s theorem since $\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))\in\operatorname{Dom}g$ for all $x\in{}^{n}$, and $\operatorname{Dom}g=\operatorname{Dom}F$. ∎
@@
For a lower semicontinuous function, we can establish that it is nonincreasing provided that its upper right Dini derivative is nonpositive [10, Theorem 1.13]. We note that $F(\phi_{t}(x))$ is lower semicontinuous, as $g$ is CCP and hence lower semicontinuous, which implies $F$ is as well, which in turn implies that the composition $t\mapsto F(\phi_{t}(x))$ is lower semicontinuous. Now, we analyse each case separately. If $F(\phi_{t}(x))=+\infty$ for all $t\geq 0$, there is nothing to prove. Suppose alternatively that it is always finite. Then we know that the map $F(\phi_{t}(x))$ is lower semicontinuous and we can compute its Dini derivative, and show that it is nonpositive. Upon noting that the constant $F^{*}$ does not affect the Dini derivative $D^{+}F(\phi_{t}(x))$, we may use the same arguments obtained in Theorem 4, to obtain,
@@
In other words, when $F(\phi_{t}(x))$ is finite, it is nonincreasing. In the case that it is infinite and then becomes finite after some time horizon, the same argument as above applies on the interval of time when $F(\phi_{t}(x))$ is finite. Due to the forward-invariance of $\operatorname{Dom}g$, $F(\phi_{t}(x))$ is either always equal to +$\infty$+ ∞, is always finite, or is infinite for a finite amount of time and then becomes finite. In other words, once $F(\phi_{t}(x))$ is finite, it can never become +$\infty$+ ∞ again. ∎
@@
IV Polyak-Łojaciewicz conditions for proximal gradient dynamics
Next, we discuss sufficient conditions for exponential convergence of the cost function of (3) under proximal gradient dynamics (5).

### v2

@@ Abstract
In this letter we study the proximal gradient dynamics. This recently-proposed continuous-time dynamics solves optimization problems whose cost functions are separable into a nonsmooth convex and a smooth component. First, we show that the cost function decreases monotonically along the trajectories of the proximal gradient dynamics. We then introduce a new condition that guarantees exponential convergence of the cost function to its optimal value, and show that this condition implies the proximal Polyak-Łojasiewicz condition. We also show that the proximal Polyak-Łojasiewicz condition guarantees exponential convergence of the cost function. Moreover, we extend these results to time-varying optimization problems, providing bounds for equilibrium tracking. Finally, we discuss applications of these findings, including the LASSO problem, certain matrix based problems and a numerical experiment on a feed-forward neural network.
@@ I Introduction
Optimization problems are central to various domains of engineering and data science. Many of these problems can be framed as minimizing a smooth cost function combined with a nonsmooth but convex regularizer that promotes structure in the solution. For example, the addition of the $\ell_{1}$ norm as a regularizer induces sparsity, a key principle in compressed sensing and sparse coding [6]. In other cases, convex constraints can be incorporated using indicator functions. A common approach for solving such problems, where the cost function is separable into a smooth, possibly nonconvex, and a nonsmooth convex component, is the proximal gradient algorithm [21]. This (discrete-time) algorithm gained prominence in sparse encoding applications [6]. A list of widely used regularizers and their associated proximal operators is available in [5, Chapter 10].
Analyzing the continuous-time versions of discrete-time optimization algorithms has garnered much interest since the seminal work in [3]. Continuous-time dynamics have proven useful across various engineering disciplines, including notable contributions from Hopfield and Tank in dynamical neuroscience [17], Kennedy and Chua in analog circuit design [20], and Brockett in systems and control theory [9]. There are several compelling reasons to investigate continuous-time version of the proximal gradient algorithm. First, these dynamics establish a connection between machine learning and dynamical neuroscience. The inclusion of a nonsmooth convex component allows for the incorporation of regularizers into neuroscientific loss landscapes, similar to those used in machine learning. Second, proximal algorithms form the cornerstone of modern optimization algorithms. In Model Predictive Control (MPC), solution methods like Sequential Quadratic Programming (SQP) and Alternating Direction Method of Multipliers (ADMM) rely on proximal gradient updates. Insights from continuous-time algorithms also inform discrete-time approaches.
@@
Related Work: The analysis of algorithms through the lens of systems theory is of much recent interest [13]. Proximal gradient algorithms have been extensively studied in the discrete-time optimization setting [21]. A key result from this analysis is the identification of the Polyak-Łojasiewicz (PL) condition, initially introduced for gradient descent in [22] and later extended to proximal methods in [19]. When the cost function satisfies this condition in discrete time, proximal gradient algorithms achive linear convergence. The authors of [19] also discuss several existing conditions in literature and their relationships with the PL condition. Notably, they propose proximal versions of the PL and Kurdyka Łojasiewicz (KL) conditions and show their equivalence.
@@
The study of optimization problems as continuous-time dynamical systems has a long history [3]. More recently, they have been used to solve primal dual dynamics [23], nonlinear programming problems [2] and saddle point problems [11]. The continuous-time proximal gradient dynamics were first proposed in [16] where it was shown that the PL condition implies the exponential convergence of the forward-backward envelope, even in the absence of strong convexity.
@@
Contributions: In this work, we analyze the convergence properties and applications of proximal gradient dynamics for optimization problems with cost functions that combine a nonsmooth convex term and a potentially nonconvex smooth term in continuous time. Our analysis is based on studying the convergence of the cost function, as in many applications, the cost is typically measurable and interpretable. Our first contribution in Theorem 2 demonstrates that the cost function decreases monotonically along the trajectories of the proximal gradient dynamics, even in a general nonconvex setting. This finding draws a parallel with the behavior of gradient descent dynamics in continuous time. We also some general properties of subgradients of convex functions, which facilitate the proof of Theorem 2. We then identify a natural condition that ensures exponential convergence in continuous time; we show that our condition implies the proximal KL condition. Next, we examine the PL condition for proximal gradient dynamics in Theorem 9 and prove that it guarantees exponential convergence of the cost function, which is an exact analog to the result in discrete time. Prior work [16] showed that, under the PL condition, the forward backward envelope of the cost function exponentially decays, but this function is not necessarily interpretable in all cases. Additionally, we derive Lipschitz-based bounds for tracking the optimal function value in parameter-varying problems under the PL condition in Theorem 12. Finally, we explore applications of these results in sparse reconstruction, neural networks, and the analysis of some matrix based problems.
@@
The letter is structured as follows: Section II covers preliminaries and formulates the optimization problem. Section III presents our main result on the convergence of the proximal dynamics. We discuss exponential stability under the proximal PL condition in Section IV, followed by applications of our results in Section V.
@@ II-A Proximal operator and the proximal gradient dynamics
The proximal gradient dynamics proposed in [16], are given by
@@ II-B Polyak-Łojaciewicz and related conditions
A differentiable function $f\colon{}^{n}\rightarrow\real$ which has a minimum at $x^{\star}\in\mathbb{R}^{n}$ is said to satisfy the PL condition if there exists a $\mu>0$ such that
@@
The PL condition is weaker than the usual assumption of convexity. A discussion of the different conditions for convergence of gradient descent in discrete time is discussed in [19]. Further [19] also proposes an extension of the PL condition for the proximal case, under the assumption of $L-$smoothness, and showed its equivalence to the KL condition.
@@
$f\colon{}^{n}\rightarrow\real$ and $g\colon{}^{n}\rightarrow\overline{\real}$ are said to satisfy the proximal PL condition if there exists a ${\mu>0}$ and ${\alpha>0}$, for all $x$, such that
@@ III Global convergence of the proximal gradient dynamics
We begin with our first main result which shows that under the proximal gradient dynamics, the cost function decreases along trajectories of the system. First, we consider the case where $g$ is finite-valued.
@@
For the optimization problem (3), let the following assumptions hold true.
@@
    For the proximal gradient dynamics (5), let $\phi_{h}\colon{}^{n}\rightarrow{}^{n}$ denote the $h$-time flow map of the dynamics. Under the assumption of continuously differentiable $f$, $\operatorname{F_{\mathrm{prox}}}$ is locally Lipschitz, and thus the flow map $\phi_{h}$ exists for sufficiently small $h\in\real$ forward in time. We consider the function $V(x)=F(x)-F^{*}$, and consider its Dini derivative with respect to time, denoting $x(t)$ by $x$ for brevity, setting ${z=\mathrm{prox}_{\alpha g}(x-\alpha\nabla f(x))}$. We will let $\mathcal{A}$ be the set of nonincreasing sequences with positive entries that converge to 00 and let $\mathcal{X}$ be the set of (vector-valued) sequences converging to $x$. For a sequence $\{x_{k}\}_{k=1}^{\infty}\in\mathcal{X}$, define $\{s_{x_{k}}\}_{k=1}^{\infty}$ to be the sequence defined by $s_{x_{k}}=\underset{s\in{}^{n}}{\textup{arg\,min}}\|s\|_{2}^{2}$ s.t. $s\in\partial g(x_{k})$ for all $k$. Then,
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D^{+}V(x)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\limsup_{h\to 0^{+}}\frac{V(\phi_{h}(x))-V(x)}{h}}$
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overset{(!)}{\leq}\nabla f(x)^{\top}(z-x)+\limsup_{h\to 0^{+}}\frac{g(\phi_{h}(x))-g(x)}{h}}$
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overset{(\star)}{\leq}\nabla f(x)^{\top}(z-x)+\limsup_{h\to 0^{+}}\frac{s_{\phi_{h}(x)}^{\top}(\phi_{h}(x)-x)}{h}}$
    (æ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overset{(\ae)}{\leq}\nabla f(x)^{\top}(z-x)+\limsup_{h\to 0^{+}}s_{\phi_{h}(x)}^{\top}(z-x)}$
    =${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overset{(\diamond)}{=}\nabla f(x)^{\top}(z-x)+\sup_{\{h_{k}\}\in\mathcal{A}}\limsup_{k\to\infty}s_{\phi_{h_{k}}(x)}^{\top}(z-x)}$
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overset{(\pi)}{\leq}\nabla f(x)^{\top}(z-x)+\sup_{\{x_{k}\}\in\mathcal{X}}\limsup_{k\to\infty}s_{x_{k}}^{\top}(z-x)}$
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overset{\varpi}{\leq}\nabla f(x)^{\top}(z-x)+\max_{{s_{x}\in\partial g(x)}}s_{x}^{\top}(z-x),}$ (10)
    where $s_{\phi_{h}(x)}$ is the minimum-norm element in $\partial g(\phi_{h}(x))$. We clarify each of the above (in)equalities. Inequality (!)(!)( ! ) holds because $D^{+}(\varphi+\theta)(t)\leq D^{+}\varphi(t)+D^{+}\theta(t)$ and $D^{+}f(x)=\nabla f(x)^{\top}(z-x)$. Inequality ()($\star$)( ⋆ ) holds because $g$ is convex and $g(x)\geq g(\phi_{h}(x))+s_{\phi_{h}(x)}^{\top}(x-\phi_{h}(x))$ for all $s_{\phi_{h}(x)}\in\partial g(\phi_{h}(x))$. Inequality $(\ae)$ holds because $x-\phi_{h}(x)=h(z-x)+\mathcal{O}(h^{2})$. Finally, inequality $(\varpi)$ is technical, but is essentially a semicontinuity argument and is proved in Lemma 16 for $s_{x}\in\partial g(x)$. By convexity of $g(x)$,
    $D^{+}V(x(t))\leq\nabla f(x)^{\top}(z-x)+g(z)-g(x).$ (11)
@@
{arxiv}
Remark 5:
In prior work [8], a similar result was proved for the special case where $g(x)$ is the indicator of a convex set. Our result generalizes to all proximal operators.
@@
Under the same assumptions as Theorem 2, the condition (9) implies the KL condition (8), and by extension also implies the PL condition (7).
{lcss}
{arxiv}
@@
    The KL condition follows upon squaring both sides and substituting (9). ∎
In Theorem 2, we show that costs along the trajectories of the proximal gradient dynamics are decreasing, in cases where $g$ is finite-valued. In many practical examples, for instance, in the case where $g$ is an indicator function, this assumption is not met. However, in Theorem 7, we will show that the cost remains non-increasing even for non-finite-valued $g$.
For the proximal gradient dynamics (5), recall that $\phi_{h}\colon{}^{n}\rightarrow{}^{n}$ denotes the $h$-time flow map of the dynamics. We consider the evolution of $F(x)=f(x)+g(x)$ along trajectories of the system by studying the evolution of $t\mapsto F(\phi_{t}(x))$ for all $x$.
@@
For a lower semicontinuous function, we can establish that it is nonincreasing provided that its upper right Dini derivative is nonpositive [14, Theorem 1.13]. We note that $F(\phi_{t}(x))$ is lower semicontinuous, as $g$ is CCP and hence lower semicontinuous, which implies $F$ is as well, which in turn implies that the composition $t\mapsto F(\phi_{t}(x))$ is lower semicontinuous. Now, we analyse each case separately. If $F(\phi_{t}(x))=+\infty$ for all $t\geq 0$, there is nothing to prove. Suppose alternatively that it is always finite. Then we know that the map $F(\phi_{t}(x))$ is lower semicontinuous and we can compute its Dini derivative, and show that it is nonpositive. Upon noting that the constant $F^{*}$ does not affect the Dini derivative $D^{+}F(\phi_{t}(x))$, we may use the same arguments obtained in Theorem 2, to obtain
@@
In other words, when $F(\phi_{t}(x))$ is finite, it is nonincreasing. In the case that it is infinite and then becomes finite after some time horizon, the same argument as above applies on the interval of time when $F(\phi_{t}(x))$ is finite. Due to the forward-invariance of $\operatorname{Dom}g$ and Lemma 15, $F(\phi_{t}(x))$ is either always equal to +$\infty$+ ∞, is always finite, or is infinite for a finite amount of time and then becomes finite. In other words, once $F(\phi_{t}(x))$ is finite, it can never become +$\infty$+ ∞. ∎
@@
IV Exponential convergence of the cost
We discuss sufficient conditions for exponential convergence of the cost in (3) under our dynamics (5).